OpenStudy (anonymous):

A tower casts a shadow 48 m long. Tom is 2 m tall and casts a shadow 8 m long while standing beside the tower. How tall is the tower?

OpenStudy (anonymous):

since the two shadows are being made at the same time when the angle of light is same, the ratio of object size to shadow size will be constant Let the size of tower be x m x 2 -- = ---- 48 8 2*48 x = ------ = 2*6 = 12 m 8 so the tower is 12 m tall